OpenStudy (lastdaywork):

Please explain the inconsistency - Given ω = (-1+i√3)/2 Find (ω^2 - ω)^2

OpenStudy (lastdaywork):

Already found the error.. XD

OpenStudy (ikram002p):

lol ok i got - 27/4

OpenStudy (lastdaywork):

Shouldn't it be (-3) ??

OpenStudy (ikram002p):

why ?

OpenStudy (lastdaywork):

ω is a complex cube root of unity ; hence ω^2 will be its conjugate.

OpenStudy (ikram002p):

ω^2=(1-3-2i√3)/4 ω^2 - ω=(-3√3)/4

OpenStudy (ikram002p):

ohh lol ok ok my bad sry

OpenStudy (lastdaywork):

ω^2=(1-3-2i√3)/4 = (-1 - i√3) / 2 ω^2 - ω = (-i√3) Earlier, I missed the 'i' and hence got confused over the negative sign which I got from the following solution - (ω^2 - ω)^2 = ω^4 + ω^2 - 2ω^3 = ω + ω^2 - 2 = -3